How many terms are there for the arithmetic sequence


28 , 25 , 22 , 19 , ⋯ , − 29 , − 32 ?


1
Expert's answer
2022-02-16T07:55:54-0500
"a_1=28, d=a_2-a_1=25-28=-3"

The n-th term of the arithmetic sequence is

"a_n=a_1+d(n-1)"

"a_n=28-3(n-1)=-32"

"3(n-1)=60"

"n-1=20"

"n=21"

There are 21 terms in the given arithmetic sequence.